# Beyond Simple Graphs: Neural Multi-Objective Routing on Multigraphs

### v4

@@ Appendix A Extended Motivation
In summary, while it is possible to work on transformed simple graphs instead of underlying multigraphs, we obtain significant advantages by not doing so. Optimizing architectures on transformed graphs constitutes another research direction that would be interesting to pursue, but is beyond the scope of this work. As shown by the occasionally poor performance of MBM on asymmetric simple graphs in Table 1, it seems unlikely that such an ”off-the-shelf” architecture would work well on much larger graphs induced by transforming multigraphs.
Appendix B Extended Related Work
B.1 Single-Objective Learning for Routing
In the non-autoregressive category, some works are Joshi et al. (2019); Fu et al. (2021); Qiu et al. (2022); Sun & Yang (2023). Typically, these methods require more sophisticated procedures than simple sampling in the decoding, such as Monte–Carlo tree search, to perform competitively, but scale better to large instances due to lower space and time complexity. Thus, an interesting recent work combining autoregressive and non-autoregressive construction to solve large instances (up to $10^{5}$ nodes for the TSP) while maintaining performance is Ye et al. (2024).
Besides constructive approaches, improvement-based methods usually obtain great results, but with longer running-times, by refining initial solutions. Some examples are Chen & Tian (2019); Lu et al. (2020); Hottung & Tierney (2020); Hudson et al. (2022).
B.2 Multi-Objective Learning for Routing
Early works treating MO routing using learning utilize multiple models, where one model is responsible for one preference between the objectives (Li et al., 2021; Wu et al., 2020). This quickly becomes infeasible as the number of models required grows exponentially with the number of objectives (Ehrgott, 2005). Pareto-set learning for multi-objective routing using a single model was proposed by Lin et al. (2022). They used the hyper-networks of Navon et al. (2021) to condition the decoder on the preference. Recent studies that feature improved results with a similar conditioning method include Fan et al. (2025) and Lu et al. (2024). An alternative approach is to use meta-learning to train a single model that can be quickly fine-tuned over a few steps to solve individual subproblems. Two representative works in this direction are Zhang et al. (2021) and Chen et al. (2023a).
Another significant recent work is that of Chen et al. (2023b), in which two diversity enhancement methods are introduced. The authors combine these methods with both the meta-learning approach and hyper-network approach to achieve improved results.
Besides tackling asymmetric problems, our work is novel in the MO routing setting as most methods utilize autoregressive construction, while one of our models combine autoregressive and non-autoregressive construction.
B.3 Routing on Multigraphs
@@ C.3 MatNet-Based Benchmark Model
The MatNet-architecture operates on bipartite graphs. It utilizes mixed-score attention layers together with fully-connected layers and normalization as in standard transformer layers (Vaswani et al., 2017). These attention layers ”mix” internally computed attention scores with external information from e.g., a distance matrix $D$. To accommodate the use of many edge features represented by distance matrices, $D_{1},\dots,D_{f}$, in our multi-objective setting, we follow the idea in Appendix A of the original paper. That is, we mix the internal attention score with all distance matrices using an MLP with $f+1$ inputs.
For the node-input to the initial MatNet-layer, in the absence of node-features we use zero-vectors for the ”row” nodes and one-hot vectors for the ”column” nodes, as in the original paper. If there are node features, e.g., demand in the MOCVRP, we use these as ”row” input while keeping the one-hot ”column” input. This allows us to augment each instance by sampling several one-hot combinations rather than scaling the instance as is done for GMS.
[…]
@@ D.2 Training Times
0505010010015015020020090909595100100EpochHV ratio (%)MOTSP TMAT200505010010015015020020096969898100100EpochMGMOTSP FLEX2-20MBMGMS-EBGMS-DH
@@
0505010010015015020020090909595100100EpochHV ratio (%)MOTSP TMAT5005050100100150150200200949496969898100100EpochMGMOTSP FLEX2-50GMS-EBGMS-DHInstance Size Switch
@@ E.2 MOTSP
Problem Description. In the multi-objective traveling salesman problem the goal is to visit all nodes/customers and return to the starting position. The Euclidean version of this problem has been widely treated in the learning-for-routing community, e.g., by Lin et al. (2022) and Fan et al. (2025). In our case, we sample $m$ distances, corresponding to the $m$ objectives, between each pair of customers. The goal is to design a tour that minimizes the sum of the distances. A tour can start from any node.
@@
Model Details. There are no node features for this problem, hence both the GREAT-based and MatNet-based encoders have the two distances as edge features, while the latter uses zero-vectors as ”row” input.
@@ E.4 MGMOTSP
E.5 MGMOTSPTW
@@
Data Generation. Distances and travel times are sampled according to the FLEXxx and FIXxx distributions. Additionally, we sample time-windows according to the ”medium” distribution of Chen et al. (2024) and Bi et al. (2024). In this setting, time-windows are sampled uniformly and independently, which results in difficult instances where there is a true trade-off between optimal paths and paths satisfying the time-window constraint. We also argue that this premise is simple for initial evaluation of our models, especially in comparison to e.g., the ”hard” instances of Chen et al. (2024), which test out-of-distribution generalization.
@@
Model Details. In the decoder, we enhance the routing context with an embedding of the current time. In the encoder, the node features in this case are the starting time and end time of the service, while the edge features are distance and travel time. Consequently, we augment the input features of GREAT so that an edge is associated with its distance, travel time and the time-window of the ending node. As usual for MatNet, the node features are used as ”row” inputs while the edge features are used in the distance matrix.
@@ Appendix F Setup for MOEAs
As a base for the implementation of NSGA-II, NSGA-III and MOEA/D, we utilized Pymoo [^2] https://pymoo.org/ (Blank & Deb, 2020) and as a base for MOGLS, we utilized GeneticLocalSearch_TSP [^3] https://github.com/kevin031060/Genetic_Local_Search_TSP (Li et al., 2021). For the former three, we limit the number of iterations to $1000$ while the latter is limited to $10000$. Additionally, local search strategies are a necessity for efficient usage of genetic algorithms in combinatorial optimization. For the MOTSP, we utilized the 2-opt heuristic of Jaszkiewicz (2002), while we utilized the problem-specific local search and chromosome decoding strategy of Lacomme et al. (2006) for the MOCVRP.
@@ Appendix H Complementary Numerical Results
H.1 Problems on Simple Graphs
Table 6 and 7 show results for the MOTSP and MOCVRP respectively. Apart from the benchmarks in the main text, we present the performance of three simple greedy heuristics on the MOTSP: Nearest Neighbor (NN), Nearest Insertion (NI) and Farthest Insertion (FI). Like LKH and OR-tools, these utilize the linear scalarization to transform the MO problem into single-objective subproblems. Additionally, we show results on Euclidean instances (EUC). For these, there are many existing models in literature and thus we replace MBM with the Conditional Neural Heuristic (CNH) of Fan et al. (2025). This model is augmented with the Euclidean augmentation of Lin et al. (2022).
@@ H.1 Problems on Simple Graphs
H.2 Multigraph Problems
We display additional results for the MGMOTSP and MGMOTSPTW in Tables 8-10. As for the MOTSP, we include NN, NI and FI for MGMOTSP. In the case of MGMOTSPTW, we also include results for LKH and OR-tools when treating the problem as the MGMOTSP. That is, we neglect the time-windows and treat the time duration of each edge exactly as the distance. We call these methods implicit LKH and implicit OR-tools, as they implicitly minimize the time-window violations by reducing the time the tour takes. Additionally, for the MGMOTSP, we show results for FLEX55 and FIX55, i.e., when there are 5 parallel edges between each node pair.
@@ H.2 Multigraph Problems
H.3 Generalization to Larger Instances
In Tables 11-13, we report more generalization results for our methods in the MOTSP and MGMOTSP cases. Notably, even when introducing significantly more edges by testing on FLEX1010 and FIX1010, our methods remain competitive zero-shot.
@@ H.3 Generalization to Larger Instances
H.4 Tri-Objective Problems
Finally, we show results when considering three objectives, rather than two as in the previous cases. In this case, we restrict ourselves to the MOTSP and MGMOTSP problems. We display the results in Tables 14 and 15. Generally, our methods seem to scale well with the number of objectives. The exception is GMS-DH on the XASY distribution for the MOTSP with 100 nodes, which seems to underperform compared to, for example, MBM and OR tools. Otherwise, both GMS-EB and GMS-DH show competitive results with a comparably short inference time. As expected, CNH performs slightly better on the Euclidean distribution compared to our methods.

### v5

@@ B.2 Node-Based Construction and Graph Transformations
In summary, while it is possible to work on transformed simple graphs instead of underlying multigraphs, we obtain significant advantages by not doing so. Optimizing architectures on transformed graphs constitutes another research direction that would be interesting to pursue, but is beyond the scope of this work. As shown by the occasionally poor performance of MBM on asymmetric simple graphs in Table 1, it seems unlikely that such an “off-the-shelf” architecture would work well on much larger graphs induced by transforming multigraphs.
B.3 Technical Novelty
Besides being the first and only neural approach to tackle multigraph routing problems, GMS is novel in several other aspects. Firstly, GMS-EB is quite novel in its edge-selection procedure during decoding. Some earlier works in the autoregressive paradigm, e.g., Zhou et al. (2025), utilize edge selection in the decoding, but not in the context of multigraphs to differentiate between edges between the same node pair. The exact edge-based multi-pointer decoder for GMS-EB, described in Appendix C.1, is also novel, being an extension of the light-weight multi-pointer decoder from Jin et al. (2023). Moreover, our edge-based decoder utilizes weighted scalarized costs in the scoring to obtain good initial tour generation, which to our knowledge is a novel addition in the multi-objective context.
Regarding GMS-DH, the combined Non-Autoregressive (NAR) and Autoregressive (AR) nature is novel in the edge-pruning and node permutation context, and specifically designed to avoid much of the scaling issues inherent with multigraphs. Previously, e.g., by Ye et al. (2024), NAR and AR model combinations have been explored for decomposing large graphs using divide-and-conquer strategies. Also note that we utilize the same encoder for both the NAR-part and AR-part of the model in the dual-decoder setup, which is atypical (Ye et al., 2024; Pan et al., 2025) and indicates that our encoder produces expressive edge embeddings. Additionally, the collaborative RL-based training of the two heads (outlined in Section 4.3) is different compared to other dual-decoder work (Bi et al., 2024) and NAR + AR work (Ye et al., 2024). It ensures all parts of the model can be trained simultaneously without access to high-quality labels. Finally, the selection-decoder, described in Appendix C.2, has been designed for our particular pruning task of choosing between parallel edges given a preference. Consequently, it has not been studies before in previous work.
@@ C.3 MatNet-Based Benchmark Model
The MatNet-architecture operates on bipartite graphs. It utilizes mixed-score attention layers together with fully-connected layers and normalization as in standard transformer layers (Vaswani et al., 2017). These attention layers “mix” internally computed attention scores with external information from e.g., a distance matrix $D$. To accommodate the use of many edge features represented by distance matrices, $D_{1},\dots,D_{f}$, in our multi-objective setting, we follow the idea in Appendix A of the original paper. That is, we mix the internal attention score with all distance matrices using an MLP with $f+1$ inputs.
@@
For the node-input to the initial MatNet-layer, in the absence of node-features we use zero-vectors for the “row” nodes and one-hot vectors for the “column” nodes, as in the original paper. If there are node features, e.g., demand in the MOCVRP, we use these as “row” input while keeping the one-hot “column” input. This allows us to augment each instance by sampling several one-hot combinations rather than scaling the instance as is done for GMS.
@@ E.2 MOTSP
Problem Description. In the multi-objective traveling salesman problem the goal is to visit all nodes/customers and return to the starting position. The Euclidean version of this problem has been widely treated in the learning-for-routing community, e.g., by Lin et al. (2022) and Fan et al. (2025a). In our case, we sample $m$ distances, corresponding to the $m$ objectives, between each pair of customers. The goal is to design a tour that minimizes the sum of the distances. A tour can start from any node.
@@
Model Details. There are no node features for this problem, hence both the GREAT-based and MatNet-based encoders have the two distances as edge features, while the latter uses zero-vectors as “row” input.
@@ E.4 MGMOTSP
E.5 MGMOCVRP
Problem Description. Like in the MOCVRP, the goal in multigraph MOCVRP is to determine an optimal route serving all customers for a truck starting at a depot with fixed capacity. Unlike the MOCVRP described above, in the MGMOCVRP we do not consider the makespan as one objective and total distance as the other. Instead, we consider two conflicting total distances, giving rise to multiple conflicting edges between each node pair.
Scalarizing this problem and correspondingly pruning the multigraph ensures that the near-optimal baseline HGS (Vidal et al., 2012) is applicable to the scalarized problem. This ensures we can accurately assess model performance.
Data Generation. We sample distances between the nodes according to the FLEXxx and FIXxx distributions. Demands are sampled as in the MOCVRP.
Model Details. The same model setup is used as for the MOCVRP, with the exception that MBM uses linear pre-pruning of multigraphs into simple graphs (see MGMOTSP for details).
Evaluation Details. For the FLEXxx distribution we utilize $(15,15)$, $(40,40)$ and $(60,60)$ as references with $20$, $50$ and $100$ nodes. For FIXxx, we always set the reference to $(N,N)$, where $N$ is the number of customers in the problem.
E.6 MGMOTSPTW
@@
Data Generation. Distances and travel times are sampled according to the FLEXxx and FIXxx distributions. Additionally, we sample time-windows according to the “medium” distribution of Chen et al. (2024) and Bi et al. (2024). In this setting, time-windows are sampled uniformly and independently, which results in difficult instances where there is a true trade-off between optimal paths and paths satisfying the time-window constraint. We also argue that this premise is simple for initial evaluation of our models, especially in comparison to e.g., the “hard” instances of Chen et al. (2024), which test out-of-distribution generalization.
@@
Model Details. In the decoder, we enhance the routing context with an embedding of the current time. In the encoder, the node features in this case are the starting time and end time of the service, while the edge features are distance and travel time. Consequently, we augment the input features of GREAT so that an edge is associated with its distance, travel time and the time-window of the ending node. As usual for MatNet, the node features are used as “row” inputs while the edge features are used in the distance matrix.
@@ Appendix F Setup for MOEAs
As a base for the implementation of NSGA-II, NSGA-III and MOEA/D, we utilized Pymoo [^2] https://pymoo.org/ (Blank and Deb, 2020) and as a base for MOGLS, we utilized GeneticLocalSearch_TSP [^3] https://github.com/kevin031060/Genetic_Local_Search_TSP (Li et al., 2021). For the former three, we limit the number of iterations to $1000$ while the latter is limited to $10000$. Additionally, local search strategies are a necessity for efficient usage of genetic algorithms in combinatorial optimization. For the MOTSP, we utilized the 2-opt heuristic of Jaszkiewicz (2002), while we utilized the problem-specific local search and chromosome decoding strategy of Lacomme et al. (2006) for the MOCVRP.
@@ Appendix H Complementary Numerical Results
H.1 Additional Neural Baselines
Here, we show more results for naive extensions of single-objective neural solvers for asymmetric problems, applied to multi-objective asymmetric problems and multigraph problems. More specifically, we implement a GREAT-Based Model (GRBM) and a GOAL-Based Model (GOBM), utilizing GREAT NB layers (Lischka et al., 2025) and GOAL mixed-attention layers (Drakulic et al., 2025) in the encoder. The rest of the framework is the same as for MBM. That is, we utilize the POMO framework for training and preference-conditioned multi-pointer networks for decoding. In the multigraph scenarios, we also pre-prune the graphs to handle edge selection and to ensure the the input-structure is compatible for GOBM.
The corresponding results are displayed in Table 6. While GRBM has a slightly lower gap compared to GMS-DH on TMAT50 and XASY50 for the MOCVRP, and GOBM has a slightly lower gap on FIX2-100 for the MGMOTSP, both GRBM and GOBM fail to perform robustly across distributions, graph sizes and problems. In contrast, GMS-DH and GMS-EB maintain a strong performance irrespective of the graph distribution and routing problem. GMS-DH performs better than GRBM and GOBM in 13 of the 16 cases displayed in Table 6, often quite significantly.
Table 6: Additional neural baselines for the MOTSP, MOCVRP, MGMOTSP and MGMOCVRP across graph sizes (50 and 100 nodes) and distributions. We also include GMS-DH for easy performance comparisons. Note that gap is with respect to the best performing method in Table 1.
TMAT50 TMAT100 XASY50 XASY100
HV Gap Time HV Gap Time HV Gap Time HV Gap Time
MOTSP GOBM 0.49 14.92% (3.6s) 0.55 13.95% (19s) 0.77 7.35% (3.6s) 0.83 7.88% (19s)
GOBM (aug) 0.50 13.06% (33s) 0.55 12.91% (2.7m) 0.79 5.16% (33s) 0.84 6.56% (2.7m)
GRBM 0.50 14.53% (5.5s) 0.56 11.53% (20s) 0.80 4.20% (5.6s) 0.84 6.54% (20s)
GRBM (aug) 0.50 14.06% (33s) 0.56 11.13% (2.9m) 0.81 3.02% (33s) 0.85 6.05% (2.9m)
TMAT50 TMAT100 XASY50 XASY100
MOCVRP GOBM 0.45 5.60% (6.2s) 0.42 6.58% (28s) 0.67 11.22% (6.2s) 0.66 16.67% (28s)
GOBM (aug) 0.45 4.29% (40s) 0.43 4.94% (3.5m) 0.70 7.23% (40s) 0.70 12.48% (3.5m)
GRBM 0.47 0.80% (10s) 0.37 18.79% (31s) 0.73 3.40% (11s) 0.77 3.64% (30s)
GRBM (aug) 0.47 0.08% (46s) 0.38 15.63% (3.9m) 0.74 1.20% (45s) 0.78 1.52% (3.9m)
FLEX2-50 FLEX2-100 FIX2-50 FIX2-100
MGMOTSP GOBM 0.86 4.91% (12s) 0.91 3.97% (46s) 0.89 3.02% (12s) 0.93 2.04% (46s)
GOBM (aug) 0.87 3.47% (1.4m) 0.92 3.08% (6.1m) 0.90 2.13% (1.4m) 0.94 1.52% (6.2m)
GRBM 0.77 14.53% (40s) 0.83 11.96% (2.7m) 0.89 3.27% (40s) 0.93 2.42% (2.7m)
GRBM (aug) 0.79 12.77% (5.1m) 0.85 9.81% (21m) 0.90 2.90% (5.1m) 0.93 2.24% (21m)
GMS-DH 0.89 1.84% (13s) 0.92 2.11% (49s) 0.91 1.73% (12s) 0.93 2.13% (52s)
FLEX2-50 FLEX2-100 FIX2-50 FIX2-100
MGMOCVRP GOBM 0.83 7.02% (13s) 0.82 8.18% (55s) 0.83 5.52% (13s) 0.86 5.60% (55s)
GOBM (aug) 0.84 5.36% (1.6m) 0.83 6.66% (7.1m) 0.84 4.15% (1.6m) 0.87 4.58% (7.1m)
GRBM 0.82 7.88% (46s) 0.80 9.51% (2.9m) 0.76 13.11% (46s) 0.80 11.91% (2.9m)
GRBM (aug) 0.83 6.54% (5.7m) 0.82 8.04% (23m) 0.80 8.23% (5.6m) 0.85 6.68% (23m)
GMS-DH 0.85 4.86% (20s) 0.85 4.94% (1.0m) 0.83 4.96% (20s) 0.88 3.58% (1.0m)
GMS-DH (aug) 0.85 3.86% (2.1m) 0.85 3.89% (8.2m) 0.84 3.96% (2.1m) 0.88 3.05% (8.1m)
H.2 Problems on Simple Graphs
Table 7 and 8 show results for the MOTSP and MOCVRP respectively. Apart from the benchmarks in the main text, we present the performance of three simple greedy heuristics on the MOTSP: Nearest Neighbor (NN), Nearest Insertion (NI) and Farthest Insertion (FI). Like LKH and OR-tools, these utilize the linear scalarization to transform the MO problem into single-objective subproblems. Additionally, we show results on Euclidean instances (EUC). For these, there are many existing models in literature and thus we replace MBM with the Conditional Neural Heuristic (CNH) of Fan et al. (2025a). This model is augmented with the Euclidean augmentation of Lin et al. (2022).
@@ H.2 Problems on Simple Graphs
H.3 Multigraph Problems
We display additional results for the MGMOTSP, MGMOCVRP and MGMOTSPTW in Tables 9-12. As for the MOTSP, we include NN, NI and FI for MGMOTSP and NN for the MGMOCVRP. In the case of MGMOTSPTW, we also include results for LKH and OR-tools when treating the problem as the MGMOTSP. That is, we neglect the time-windows and treat the time duration of each edge exactly as the distance. We call these methods implicit LKH and implicit OR-tools, as they implicitly minimize the time-window violations by reducing the time the tour takes. Additionally, for the MGMOTSP, we show results for FLEX55 and FIX55, i.e., when there are 5 parallel edges between each node pair.
@@ H.3 Multigraph Problems
Table 11: MGMOCVRP on FLEXxx and FIXxx with different instances sizes (20, 50 and 100).
FLEX2-20 FLEX2-50 FLEX2-100
HV Gap Time HV Gap Time HV Gap Time
HGS 0.76 0.00% (2.3h) 0.89 0.00% (15h) 0.89 0.00% (75h)
OR Tools 0.71 6.30% (1.3h) 0.84 5.93% (1.5h) 0.84 5.97% (1.9h)
NN 0.55 28.04% (4.5s) 0.72 18.73% (18s) 0.73 17.52% (1.0m)
MOGLS 0.54 29.34% (4.4h) 0.46 48.56% (49h) 0.24 73.18% (122h)
MOEAD 0.65 14.14% (5.9h) 0.68 23.83% (36h) 0.56 37.47% (218h)
NSGA-II 0.69 9.78% (5.7h) 0.71 20.05% (37h) 0.56 36.54% (219h)
NSGA-III 0.68 11.02% (5.8h) 0.70 21.32% (36h) 0.55 38.08% (224h)
MBM 0.72 5.68% (5.7s) 0.83 6.07% (16s) 0.81 9.38% (54s)
MBM (aug) 0.73 4.34% (37s) 0.84 5.23% (2.0m) 0.82 8.30% (6.9m)
GMS-EB 0.71 6.06% (5.6s) 0.85 4.12% (1.2m) 0.85 4.08% (10m)
GMS-EB (aug) 0.73 4.36% (43s) 0.86 3.55% (9.0m) 0.86 3.66% (1.4h)
GMS-DH 0.70 7.83% (4.3s) 0.85 4.86% (20s) 0.85 4.94% (1.0m)
GMS-DH (aug) 0.72 4.94% (24s) 0.85 3.86% (2.1m) 0.85 3.89% (8.2m)
FIX2-20 FIX2-50 FIX2-100
HGS 0.77 0.00% (2.5h) 0.87 0.00% (15h) 0.91 0.00% (74h)
OR Tools 0.74 4.68% (1.3h) 0.83 5.14% (1.5h) 0.88 3.78% (1.9h)
NN 0.61 20.78% (4.5s) 0.74 15.38% (18s) 0.81 10.74% (1.0m)
MOGLS 0.54 30.14% (4.4h) 0.46 46.95% (49h) 0.36 60.78% (121h)
MOEAD 0.62 19.81% (6.3h) 0.62 28.76% (36h) 0.60 34.12% (217h)
NSGA-II 0.66 15.14% (6.1h) 0.65 25.16% (36h) 0.61 33.28% (224h)
NSGA-III 0.65 16.03% (6.2h) 0.64 26.86% (38h) 0.59 35.19% (220h)
MBM 0.73 5.87% (6.0s) 0.84 4.14% (16s) 0.87 5.00% (54s)
MBM (aug) 0.74 3.69% (38s) 0.84 3.39% (1.9m) 0.87 4.32% (6.9m)
GMS-EB 0.74 4.67% (5.5s) 0.84 3.47% (1.2m) 0.89 2.81% (9.9m)
GMS-EB (aug) 0.74 3.71% (44s) 0.85 2.71% (11m) 0.89 2.38% (1.4h)
GMS-DH 0.73 5.95% (4.3s) 0.83 4.96% (20s) 0.88 3.58% (1.0m)
GMS-DH (aug) 0.74 3.87% (24s) 0.84 3.96% (2.1m) 0.88 3.05% (8.1m)
@@
H.4 Generalization to Larger Instances
In Tables 13-15, we report more generalization results for our methods in the MOTSP and MGMOTSP cases. Notably, even when introducing significantly more edges by testing on FLEX1010 and FIX1010, our methods remain competitive zero-shot. The performance of GMS-DH degrades somewhat for FIX1010-100, which we attribute to the growing action space of the selection-decoder.
@@ H.4 Generalization to Larger Instances
H.5 Tri-Objective Problems
Finally, we show results when considering three objectives, rather than two as in the previous cases. In this case, we restrict ourselves to the MOTSP and MGMOTSP problems. We display the results in Tables 16 and 17. Generally, our methods seem to scale well with the number of objectives. The exception is GMS-DH on the XASY distribution for the MOTSP with 100 nodes, which seems to underperform compared to, for example, MBM and OR tools. Otherwise, both GMS-EB and GMS-DH show competitive results with a comparably short inference time. As expected, CNH performs slightly better on the Euclidean distribution compared to our methods.